### v2

# Mora: Enabling Generalist Video Generation via A Multi-Agent Framework

Abstract
Sora is the first large-scale generalist video generation model that garnered significant attention across society. Since its launch by OpenAI in February 2024, no other video generation models have paralleled Sora’s performance or its capacity to support a broad spectrum of video generation tasks. Additionally, there are only a few fully published video generation models, with the majority being closed-source. To address this gap, this paper proposes a new multi-agent framework Mora, which incorporates several advanced visual AI agents to replicate generalist video generation demonstrated by Sora. In particular, Mora can utilize multiple visual agents and successfully mimic Sora’s video generation capabilities in various tasks, such as (1) text-to-video generation, (2) text-conditional image-to-video generation, (3) extend generated videos, (4) video-to-video editing, (5) connect videos and (6) simulate digital worlds. Our extensive experimental results show that Mora achieves performance that is proximate to that of Sora in various tasks. However, there exists an obvious performance gap between our work and Sora when assessed holistically. In summary, we hope this project can guide the future trajectory of video generation through collaborative AI agents.
Figure 1: Samples for text-to-video generation of Mora. Our approach can generate high-resolution, temporally consistent videos from text prompts. The samples shown are 1024$\times$576 resolution over 12 seconds duration at 75 frames in total.
1 Introduction
Since the release of ChatGPT in November 2022, the advent of generative AI technologies has marked a significant transformation, reshaping interactions and integrating deeply into various facets of daily life and industry OpenAI (2023a, b). In the realm of visual AI, the focus has predominantly been on image generation with models such as Midjourney Midjourney (2023), Stable Diffusion Rombach et al. (2022a), and DALL-E 3 Betker et al. (2023) leading the way. In contrast, video generation has remained unexplored, especially when compared to the progress seen in language and image synthesis technologies. Although recent notable video generation models like Pika pik and Gen-2 Gen have demonstrated their capacity to produce diverse and high-quality videos, they have been limited by their ability to create longer-duration videos exceeding 10 seconds. This scenario has seen a revolutionary shift with the introduction of OpenAI’s Sora OpenAI (2024a), marking a new era in video generation capabilities.
Sora, a pioneering text-to-video generative model introduced by OpenAI in February 2024, distinguishes itself by its ability to convert text prompts into detailed videos, demonstrating remarkable potential in replicating physical world dynamics. This model transcends its predecessors by generating videos up to a minute long, closely aligning with the provided text descriptions OpenAI (2024a). Beyond mere text-to-video generation, Sora excels in a range of video tasks as a generalist model, including editing, connecting, and extending footage in ways previously unachieved. In addition, the generated content is known for its multi-view perspectives and fidelity to user instructions, positioning it uniquely amongst video generation models. As we look to the future, the implications of Sora and similar advanced video generation technologies are poised to make profound contributions across various sectors, including but not limited to filmmaking Zhu et al. (2023); Zhuang et al. (2024), robotics Kapelyukh et al. (2023); Liu et al. (2022), and healthcare Awasthi et al. (2023).
Despite its innovative contributions, Sora’s closed-source nature, similar to most video generation models, presents significant challenges for the academic community. The inaccessibility hinders researchers’ ability to replicate or extend Sora’s capabilities. There is a growing trend of attempts to reverse-engineer Sora, with some studies, such as Liu et al. (2024a), proposing potential techniques that might be employed within Sora, including diffusion transformers and spatial patch strategies Sohl-Dickstein et al. (2015); Ho et al. (2020); Bao et al. (2023); Peebles and Xie (2023). Despite these efforts, achieving the same level of performance and adaptability as Sora proves to be an immense challenge. The lack of access to comparable computational power and extensive training datasets further complicates these endeavors.
Table 1: Task comparison between Sora, Mora and other existing models.
Tasks Example Sora Mora Others
Text-to-video Generation
Girdhar et al. (2023); Wang et al. (2023a); Chen et al. (2024); Ma et al. (2024a)
Text-conditional Image-to-Video Generation
Blattmann et al. (2023); pik ; Gen
Extend Generated Videos
Video-to-Video Editing
Molad et al. (2023); Liew et al. (2023); Ceylan et al. (2023)
Connect Videos
Chen et al. (2023a)
Simulate Digital Worlds
To address the limitations of current video generation models, we explore the potential of multi-agent collaboration Hong et al. (2023); Wu et al. (2023a) in accomplishing generalist video generation tasks. We introduce a multi-agent framework, referred to as Mora, that leverages various advanced large models to enable text-to-video capabilities similar to Sora. Specifically, we decompose video generation into several subtasks, with each subtask assigned to a dedicated agent: (1) enhancing prompts provided by the user, (2) generating an image from an input text prompt, (3) editing or refining images based on the enhanced conditioning provided by the text, (4) generating a video from the generated image, and (5) connecting two videos. By automatically organizing agents to loop and permute through these subtasks, Mora can complete a wide array of video generation tasks through a flexible pipeline, thereby meeting the diverse needs of users.
Intuitively, equipping the model with both a starting image and text simplifies the video generation process, as it primarily needs to extrapolate the future progression of the image. This method stands in contrast to direct end-to-end text-to-video approaches Singer et al. (2023); Ho et al. (2022a); Chen et al. (2023b); Wang et al. (2023a); Chen et al. (2024). Our multi-agent collaboration framework distinctively produces an intermediate image or video during inference, enabling the preservation of the visual diversity, style, and quality inherent in the text-to-image model (see examples in Figure 1). This process even facilitates editing capabilities. By effectively coordinating the efforts of text-to-image, image-to-image, image-to-video and video-to-video agents, Mora can adeptly conduct a broad spectrum of video generation tasks while offering superior editing flexibility and visual fidelity, rivaling the performance of established models like Sora. A detailed comparison of tasks between Mora and Sora is presented in Table 1. This comparison demonstrates that, through the collaboration of multiple agents, Mora is capable of accomplishing the video-related tasks that Sora can undertake. This comparison highlights Mora’s adaptability and proficiency in addressing a multitude of video generation challenges.
To comprehensively assess the efficacy of Mora, we use basic metrics in publicly available video generation benchmark Vbench Huang et al. (2023) and self-defined metrics for six tasks, including text-to-video generation, text-conditional image-to-video generation, extending generated videos, video-to-video editing, connecting videos, and simulating digital worlds. Notably, Mora achieves superior performance in the text-to-video generation task than existing open-sourced models, ranking second only to Sora. In the other tasks, Mora also delivers competitive results, underscoring the versatility and general capabilities of our framework.
We summarize our contributions as follows:
  • In this paper, we introduce Mora, a groundbreaking meta-programming framework crafted to enhance multi-agent collaboration. This framework stands out for its structured yet adaptable system of agents, paired with an intuitive interface for the configuration of components and task pipelines. These features position Mora as a prime instrument for pushing forward the boundaries of generalist video generation tasks.
  • Our research reveals that the quality of video generation can be notably improved by leveraging the automated cooperation of multiple agents, including text-to-image, image-to-image, image-to-video, and video-to-video agents. This collaborative process starts with generating an image from text, followed by using both the generated image and the input text to produce a video. The process concludes with further refinement, extension, connection, and editing of the video.
  • Mora stands out for its exceptional performance across six video-related tasks, surpassing existing open-sourced models. This impressive achievement underlines the effectiveness of Mora, showcasing its potential as a versatile framework tailored for generalist video generation. The comprehensive results not only affirm the capabilities of Mora but also position it as a groundbreaking tool in the realm of video generation, promising significant advancements in how video content is created and utilized.
2 Related Work
2.1 Text-to-Video Generation
Generating videos based on textual descriptions has been long discussed. While early efforts in the field were primarily rooted in GANs Wang et al. (2020); Chu et al. (2020) and VQ-VAE Yan et al. (2021), recent breakthroughs in generative video models, driven by foundational work in transformer-based architectures and diffusion models, have advanced academic research. Auto-regressive transformers are early leveraged in video generation Wu et al. (2022); Hong et al. (2022a); Kondratyuk et al. (2023). These models are designed to generate video sequences in a frame-by-frame manner, predicting each new frame based on the previously generated frames. Parallelly, the adaptation of masked language models He et al. (2021) for visual contexts, as demonstrated by Gupta et al. (2022); Villegas et al. (2022); Yu et al. (2023); Gupta et al. (2023), underscores the versatility of transformers in video generation. The recently-proposed VideoPoet Kondratyuk et al. (2023) leverages an auto-regressive language model and can multitask on a variety of video-centric inputs and outputs.
In another line, large-scale diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020) show competitive performance in video generation Ho et al. (2022a); Singer et al. (2022); Khachatryan et al. (2023); Wu et al. (2023b); Du et al. (2024). By learning to gradually denoise a sample from a normal distribution, diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020) implement an iterative refinement process for video synthesis. Initially developed for image generation Rombach et al. (2022b); Podell et al. (2023), they have been adapted and extended to handle the complexities of video data. This adaptation began with extending image generation principles to video Ho et al. (2022b, a); Singer et al. (2022), by using a 3D U-Net structure instead of conventional image diffusion U-Net. In the follow-up, latent diffusion models (LDMs) Rombach et al. (2022a) are integrated into video generation Zhou et al. (2022); Chen et al. (2023b); Wang et al. (2023a); Chen et al. (2024), showcasing enhanced capabilities to capture the nuanced dynamics of video content. For instance, Stable Video Diffusion Blattmann et al. (2023) can conduct multi-view synthesis from a single image while Emu Video Girdhar et al. (2023) uses just two diffusion models to generate higher-resolution videos.
Researchers have delved into the potential of diffusion models for a variety of video manipulation tasks. Notably, Dreamix Molad et al. (2023) and MagicEdit Liew et al. (2023) have been introduced for general video editing, utilizing large-scale video-text datasets. Conversely, other models employ pre-trained models for video editing tasks in a zero-shot manner Ceylan et al. (2023); Couairon et al. (2023); Yang et al. (2023); Chai et al. (2023). SEINE Chen et al. (2023a) is specially designed for generative transition between scenes and video prediction. While these above-mentioned end-to-end models exhibit remarkable proficiency in specific areas, they encounter limitations in broadening their capabilities to encompass a wider range of video tasks, particularly those requiring varied types of inputs and the generation of longer-duration videos. The introduction of diffusion transformers Peebles and Xie (2023); Bao et al. (2023); Ma et al. (2024b) further revolutionized video generation, culminating in advanced solutions like Latte Ma et al. (2024a) and Sora OpenAI (2024a). Sora’s ability to produce minute-long videos of high visual quality that faithfully follow human instructions heralds a new era in video generation, promising unprecedented opportunities for creativity and expression in digital media.
2.2 AI Agents
Large models have enabled agents to excel across a broad spectrum of applications, showcasing their versatility and effectiveness. They have greatly advanced collaborative multi-agent structures for multimodal tasks in areas such as scientific research Tang et al. (2024), software development Hong et al. (2023); Qian et al. (2023) and society simulation Park et al. (2023). Compared to individual agents, the collaboration of multiple autonomous agents, each equipped with unique strategies and behaviors and engaged in communication with one another, can tackle more dynamic and complex tasks Guo et al. (2024). Through a cooperative agent framework known as role-playing, CAMEL Li et al. (2024) enables agents to collaborate and solve complex tasks effectively. Park et al. Park et al. (2023) designed a community of 25 generative agents capable of planning, communicating, and forming connections. Liang et al. Liang et al. (2023) have explored the use of multi-agent debates for translation and arithmetic problems, encouraging divergent thinking in large language models. Hong et al. Hong et al. (2023) introduced MetaGPT, which utilizes an assembly line paradigm to assign diverse roles to various agents. In this way, complex tasks can be broken down into subtasks, which makes it easy for many agents working together to complete. Xu et al. Xu et al. (2023) used a multi-agent collaboration strategy to simulate the academic peer review process. AutoGen Wu et al. (2023a) is a generic programming framework which can be used to implement diverse multi-agent applications across different domains, using a variety of agents and conversation patterns.
Motivated by existing works, we extend the principle of collaborative agents to complete vision tasks. By integrating and coordinating agents of various roles such as text-to-image and image-to-video agents, we manage to accomplish multiple video-related tasks in a modular and extensible approach.
3 Mora: A Multi-Agent Framework for Video Generation
As illustrated in Figure 2, Mora aims to provide a framework leveraging advanced AI agents to realize Text-to-Video generation. Sec 3.1 provides the agent definitions and example pipelines enabled by this framework. Sec 3.2 introduces the implementation detail of each agent.
3.1 Agent-based Video Generation
Definition and Specialization of Agents.   The definition of agents enables flexibility in the breakdown of complex work into smaller and more specific tasks. Solving different video generation tasks often requires the collaboration of agents with diverse abilities, each contributing specialized outputs. In our framework, we have 5 basic roles: prompt selection and generation agent, text-to-image generation agent, image-to-image generation agent, and image-to-video generation agent and video-to-video agent.
Figure 2: Illustration of how to use Mora to conduct video-related tasks.
  • Prompt Selection and Generation Agent: Prior to the commencement of the initial image generation, textual prompts undergo a rigorous processing and optimization phase. This critical agent can employ large language models like GPT-4, Llama Achiam et al. (2023); Touvron et al. (2023). It is designed to meticulously analyze the text, extracting pivotal information and actions delineated within, thereby significantly enhancing the relevance and quality of the resultant images. This step ensures that the textual descriptions are thoroughly prepared for an efficient and effective translation into visual representations.
  • Text-to-Image Generation Agent: The text-to-image model Rombach et al. (2022b); Podell et al. (2023) stands at the forefront of translating these enriched textual descriptions into high-quality initial images. Its core functionality revolves around a deep understanding and visualization of complex textual inputs, enabling it to craft detailed and accurate visual counterparts to the provided textual descriptions.
  • Image-to-Image Generation Agent: This agent Brooks et al. (2023) works to modify a given source image in response to specific textual instructions. The core of its functionality lies in its ability to interpret detailed textual prompts with high accuracy and subsequently apply these insights to manipulate the source image accordingly. This involves a detailed recognition of the text’s intent, translating these instructions into visual modifications that can range from subtle alterations to transformative changes. The agent leverages a pre-trained model to bridge the gap between textual description and visual representation, enabling seamless integration of new elements, adjustment of visual styles, or alteration of compositional aspects within the image.
  • Image-to-Video Generation Agent  Blattmann et al. (2023): Following the creation of the initial image, the Video Generation Model is responsible for transitioning the static frame into a vibrant video sequence. This component delves into the analysis of both the content and style of the initial image, serving as the foundation for generating subsequent frames. These frames are meticulously crafted to ensure a seamless narrative flow, resulting in a coherent video that upholds temporal stability and visual consistency throughout. This process highlights the model’s capability to not only understand and replicate the initial image but also to anticipate and execute logical progressions in the scene.
  • Video Connection Agent: Utilizing the Video-to-Video Agent, we create seamless transition videos based on two input videos provided by users. This advanced agent selectively leverages key frames from each input video to ensure a smooth and visually consistent transition between them. It is designed with the capability to accurately identify the common elements and styles across the two videos, thus ensuring a coherent and visually appealing output. This method not only improves the seamless flow between different video segments but also retains the distinct styles of each segment.
Every agent is responsible for the specific input and output. These results can be utilized for different designed tasks.
Approaches.   By setting the agents’ roles and operational skills, we can define basic workflows for different tasks. We design six text-to-video generation tasks: (1) Text-to-video generation, (2) Text-conditional image-to-video generation, (3) Extend generated videos, (4) Video-to-video editing, (5) Connect Videos, and (6) Simulate digital worlds, which are described below.
  • Task 1: Text-to-Video Generation: This task harnesses a detailed textual prompt from the user as the foundation for video creation. The prompt must meticulously detail the envisioned scene. Utilizing this prompt, the Text-to-Image agent utilizes the text, distilling themes and visual details to craft an initial frame. Building upon this foundation, the Image-to-Video component methodically generates a sequence of images. This sequence dynamically evolves to embody the prompt’s described actions or scenes, and each video is derived from the last frame from the previous video, thereby achieving a seamless transition throughout the video.
  • Task 2: Text-conditional Image-to-Video Generation: Task 2 mirrors the operational pipeline of Task 1, with a key distinction. Unlike Task 1 with only texts as inputs, Task 2 integrates both a textual prompt and an initial image into the Text-to-Image agent’s input. This dual-input approach enriches the content generation process, enabling a more nuanced interpretation of the user’s vision.
  • Task 3: Extend Generated Videos: This task focuses on extending the narrative of an existing video sequence. By taking the last frame of an input video as the starting point, the video generation agent crafts a series of new, coherent frames that continue the story. This approach allows for the seamless expansion of video content, creating longer narratives that maintain the consistency and flow of the original sequence.
  • Task 4: Video-to-Video Editing: Task 4 introduces a sophisticated editing capability, leveraging both the Image-to-Image and Image-to-Video agents. The process begins with the Image-to-Image agent, which takes the first frame of an input video and applies edits based on the user’s prompt, achieving the desired modifications. This edited frame then serves as the initial image for the Image-to-Video agent, which generates a new video sequence that reflects the requested obvious or subtle changes, offering a powerful tool for dynamic video editing.
  • Task 5: Connect Videos: The Image-to-Video agent leverages the final frame of the first input video and the initial frame of the second input video to create a seamless transition, producing a new video that smoothly connects the two original videos.
  • Task 6: Simulating Digital Worlds: This task specializes in the whole style changing for video sequences set in digitally styled worlds. By appending the phrase "In digital world style" to the edit prompt, the user instructs the Image-to-Video agent to craft a sequence that embodies the aesthetics and dynamics of a digital realm or utilize the Image-to-Image agent to transfer the real image to digital style. This task pushes the boundaries of video generation, enabling the creation of immersive digital environments that offer a unique visual experience.
3.2 Implementation Detail of Agents
Prompt Selection and Generation.  Currently, GPT-4  Achiam et al. (2023) stands as the most advanced generative model available. By harnessing the capabilities of GPT-4, we are able to generate and meticulously select high-quality prompts. These prompts are detailed and rich in information, facilitating the Text-to-Image generation process by providing the agent with comprehensive guidance.
Text-to-Image Generation.  We utilize the pretrained large text-to-image model to generate a high-quality and representative first image.
The Stable Diffusion XL (SDXL) Podell et al. (2023) is utilized for the first implementation. It introduces a significant evolution in the architecture and methodology of latent diffusion models Rombach et al. (2022b); Meng et al. (2022) for text-to-image synthesis, setting a new benchmark in the field. At the core of its architecture is an enlarged UNet backbone Ronneberger et al. (2015) that is three times larger than those used in previous versions of Stable Diffusion 2 Rombach et al. (2022b). This expansion is principally achieved through an increased number of attention blocks and a broader cross-attention context, facilitated by integrating a dual text encoder system. The first encoder is based on OpenCLIP Ilharco et al. (2021) ViT-bigG Cherti et al. (2023); Radford et al. (2021); Schuhmann et al. (2022), while the second utilizes CLIP ViT-L, allowing for a richer, more nuanced interpretation of textual inputs by concatenating the outputs of these encoders. This architectural innovation is complemented by the introduction of several novel conditioning schemes that do not require external supervision, enhancing the model’s flexibility and capability to generate images across multiple aspect ratios. Moreover, SDXL features a refinement model that employs a post-hoc image-to-image transformation to elevate the visual quality of the generated images. This refinement process utilizes a noising-denoising technique, further polishing the output images without compromising the efficiency or speed of the generation process.
Image-to-Image Generation.  Our initial framework realize uitilze InstructPix2Pix as Image-to-Image generation agent. InstructPix2Pix  Brooks et al. (2023) are intricately designed to enable effective image editing from natural language instructions. At its core, the system integrates the expansive knowledge of two pre-trained models: GPT-3  Brown et al. (2020) for generating editing instructions and edited captions from textual descriptions, and Stable Diffusion  Rombach et al. (2022a) for transforming these text-based inputs into visual outputs. This ingenious approach begins with fine-tuning GPT-3 on a curated dataset of image captions and corresponding edit instructions, resulting in a model that can creatively suggest plausible edits and generate modified captions. Following this, the Stable Diffusion model, augmented with the Prompt-to-Prompt technique, generates pairs of images (before and after the edit) based on the captions produced by GPT-3. The conditional diffusion model at the heart of InstructPix2Pix is then trained on this generated dataset. InstructPix2Pix directly utilizes the text instructions and input image to perform the edit in a single forward pass. This efficiency is further enhanced by employing classifier-free guidance for both the image and instruction conditionings, allowing the model to balance fidelity to the original image with adherence to the editing instructions.
Image-to-Video Generation.  In the Text-to-Video generation agent, video generation agents play an important role in ensuring video quality and consistency.
Our first implementation utilizes the state-of-the-art video generation model Stable Video Diffusion to generate video. The Stable Video Diffusion (SVD)  Blattmann et al. (2023) architecture introduces a cutting-edge approach to generating high-resolution videos by leveraging the strengths of LDMs Stable Diffusion v2.1  Rombach et al. (2022a), originally developed for image synthesis, and extending their capabilities to handle the temporal complexities inherent in video content. At its core, the SVD model follows a three-stage training regime that begins with text-to-image pertaining, where the model learns robust visual representations from a diverse set of images. This foundation allows the model to understand and generate complex visual patterns and textures. In the second stage, video pretraining, the model is exposed to large amounts of video data, enabling it to learn temporal dynamics and motion patterns by incorporating temporal convolution and attention layers alongside its spatial counterparts. This training is conducted on a systematically curated dataset, ensuring the model learns from high-quality and relevant video content. The final stage, high-quality video finetuning, focuses on refining the model’s ability to generate videos with increased resolution and fidelity, using a smaller but higher-quality dataset. This hierarchical training strategy, complemented by a novel data curation process, allows SVD to excel in producing state-of-the-art text-to-video and image-to-video synthesis with remarkable detail, realism, and coherence over time.
Connect Videos.  For the video connection task, we utilize SEINE Chen et al. (2023a) to connect videos. SEINE is constructed upon a pre-trained diffusion-based T2V model, LaVie Wang et al. (2023a) agent. SEINE centered around a random-mask video diffusion model that generates transitions based on textual descriptions. By integrating images of different scenes with text-based control, SEINE produces transition videos that maintain coherence and visual quality. Additionally, the model can be extended for tasks such as image-to-video animation and autoregressive video prediction.
4 Experiments
4.1 Setup
Baseline.  In the text-to-video generation, existing open-sourced works that show competitive performance are employed as the baseline models, including Videocrafter1 Chen et al. (2023c), Show-1 Zhang et al. (2023a), Pika pik , Gen-2 Gen , ModelScope Mod , LaVie-Interpolation, LaVie Wang et al. (2023b) and CogVideo Hong et al. (2022b). In the other five tasks, we compare Mora with Sora.
Basic Metrics.  For text-to-video generation, we employed several metrics from Vbench Huang et al. (2023) for evaluation from two aspects: video quality and video condition consistency.
For video quality measurement, we use six metrics. ❶ Object Consistency, computed by the DINO Caron et al. (2021) feature similarity across frames to assess whether object appearance remains consistent throughout the whole video; ❷ Background Consistency, calculated by CLIP Radford et al. (2021) feature similarity across frames; ❸ Motion Smoothness, which utilizes the motion priors in the video frame interpolation model AMT Li et al. (2023) to evaluate the smoothness of generated motions; ❹ Aesthetic Score, obtained by using the LAION aesthetic predictor lio on each video frame to evaluate the artistic and beauty value perceived by humans, ❺ Dynamic Degree, computed by employing RAFT Teed and Deng (2020) to estimate the degree of dynamics in synthesized videos; ❻ Imaging Quality, calculated by using MUSIQ Ke et al. (2021) image quality predictor trained on SPAQ Fang et al. (2020) dataset.
For measuring video condition consistency, we use two metrics. ❶ Temporal Style, which is determined by utilizing ViCLIP Wang et al. (2023c) to compute the similarity between video features and temporal style description features, thereby reflecting the consistency of the temporal style; ❷ Appearance Style, by calculating the feature similarity between synthesized frames and the input prompt using CLIP Radford et al. (2021), to gauge the consistency of appearance style.
Self-defined Metrics.  For evaluating other tasks, we also define four metrics.
❶ Video-Text Integration $VideoTI$, devised to enhance the quantitative evaluation of the model’s fidelity to textual instructions. It employs LLaVA Liu et al. (2024b) to transfer input image into textual descriptors $T_{i}$ and Video-Llama Zhang et al. (2023b) to transfer videos generated by the model into textual $T_{v}$. The textual representation of the image is prepended with the original instructional text, forming an augmented textual input $T_{mix}$. Both the newly formed text and the video-generated text will be input to BERT Devlin et al. (2018). The embeddings obtained are analyzed for semantic similarity through the computation of cosine similarity, providing a quantitative measurement of the model’s adherence to the given instructions and image.
$VideoTI=cosine(embed_{mix},embed_{v})$ (1)
where $embed_{mix}$ represents the embedding for $T_{mix}$ and $embed_{v}$ for $T_{v}$.
❷ Temporal Consistency $TCON$, designed to measure the coherence between an original video and its extended version, provides a vital tool for assessing the integrity of extended video content. For each input-output video pair, we employ ViCLIP Wang et al. (2023c) video encoder to extract their feature vectors. We then compute cosine similarity to get the score.
$TCON=cosine(V_{input},V_{output})$ (2)
❸ Temporal coherence $Tmean$, by quantifying the correlation between the intermediate generated video and the input videos based on $TCON$.
$Tmean=(TCON_{front}+TCON_{beh})/2$ (3)
where $TCON_{front}$ measures the correlation between the intermediate video and the preceding video in the time series while $TCON_{beh}$ assesses the correlation with the subsequent video. The average of these scores provides an aggregate measure of temporal coherence across the video sequence.
❹ Video Length, to evaluate the models’ efficiency in producing video content, specifically focusing on the maximum duration, measured in seconds.
Table 2: Comparative analysis of text-to-video generation performance between Mora and various other models. The Others category scores are derived from the Hugging Face leaderboard. For Our Mora, the evaluation is based on prompts generated by GPT-4, categorized into three types based on the number of moving objects in the videos: Type I (single object in motion), Type II (two to three objects in motion), and Type III (more than three objects in motion). Differences in the input prompt may account for the superiority of Mora inType II prompt’s scores over those of Sora in the relevant evaluations.
Model Video Quality Object Consistency Background Consistency Motion Smoothness Aesthetic Quality Dynamic Dgree Imaging Quality Temporal Style Video Length(s)
Others
Sora 0.797 0.95 0.96 1.00 0.60 0.69 0.58 0.35 60
VideoCrafter1 0.778 0.95 0.98 0.95 0.63 0.55 0.61 0.26 2
ModelScope 0.758 0.89 0.95 0.95 0.52 0.66 0.58 0.25 2
Show-1 0.751 0.95 0.98 0.98 0.57 0.44 0.59 0.25 3
Pika 0.741 0.96 0.96 0.99 0.63 0.37 0.54 0.24 3
LaVie-Interpolation 0.741 0.92 0.97 0.97 0.54 0.46 0.59 0.26 10
Gen-2 0.733 0.97 0.97 0.99 0.66 0.18 0.63 0.24 4
LaVie 0.746 0.91 0.97 0.96 0.54 0.49 0.61 0.26 3
CogVideo 0.673 0.92 0.95 0.96 0.38 0.42 0.41 0.07 4
Our Mora
Type I 0.782 0.96 0.97 0.99 0.60 0.60 0.57 0.26 12
Type II 0.810 0.94 0.95 0.99 0.57 0.80 0.61 0.26 12
Type III 0.795 0.94 0.93 0.99 0.55 0.80 0.56 0.26 12
Mora 0.792 0.95 0.95 0.99 0.57 0.70 0.59 0.26 12
Implementation Details.  For text-to-video generation, we follow the style of text prompts provided in the official Sora technical report OpenAI (2024b). Subsequently, we employ GPT-4 OpenAI (2023b) to produce more text under a few-shot setting. GPT-4 is also utilized to generate the same number of texts in a zero-shot setting. All generated text prompts are then input into text-to-video models for generating videos. For comparison with Sora, we utilize videos featured on its official website and technical report.
All experiments are conducted on two TESLA A100 GPUs, equipped with a substantial 2$\times$80GB of VRAM. The central processing was handled by 4×AMD EPYC 7552 48-Core Processors. Memory allocation was set at 320GB. The software environment was standardized on PyTorch version 2.0.2 and CUDA 12.2 for video generation and PyTorch version 1.10.2 and CUDA 11.6 for video evaluation.
4.2 Results
Text-to-Video Generation.  The quantitative results are detailed in Table 2. Mora showcases commendable performance across all metrics, making it highly comparable to the top-performing model, Sora, and surpassing the capabilities of other competitors. Specifically, Mora achieved a Video Quality score of 0.792, which closely follows Sora’s leading score of 0.797 and surpasses the current best open-source model like VideoCrafter1. In terms of Object Consistency, Mora scored 0.95, equaling Sora and demonstrating superior consistency in maintaining object identities throughout the videos. For Background Consistency and Motion Smoothness, Mora achieved scores of 0.95 and 0.99, respectively, indicating high fidelity in background stability and fluidity of motion within generated videos. Although Sora achieved 0.96 slightly outperforms Mora in Background Consistency, the margin is minimal. The Aesthetic Quality metric, which assesses the overall visual appeal of the videos, saw Mora scoring 0.57. This score, while not the highest, reflects a competitive stance against other models, with Sora scoring slightly higher at 0.60. Nevertheless, Mora’s performance in Dynamic Degree and Imaging Quality, with scores of 0.70 and 0.59, showcases its strength in generating dynamic, visually compelling content that surpasses all other models. As for Temporal Style, Mora scored 0.26, indicating its robust capability in addressing the temporal aspects of video generation. Although this performance signifies a commendable proficiency, it also highlights a considerable gap between our model and Sora, the leader in this category with a score of 0.35.
In Figure 1, the visual fidelity of Mora’s text-to-video generation is compelling, manifesting high-resolution imagery with acute attention to detail as articulated in the accompanying textual descriptions. The vivid portrayal of scenes, from the liftoff of a rocket to the dynamic coral ecosystem and the urban skateboarding vignette, underscores the system’s adeptness in capturing and translating the essence of the described activities and environments into visually coherent sequences. Notably, the images exude a temporal consistency that speaks to Mora’s nuanced understanding of narrative progression, an essential quality in video synthesis from textual prompts.
Text-conditional Image-to-Video Generation.  Analyzing Table 3, we discern a notable demonstration of Mora’s capabilities in text-conditional image-to-video generation, closely trailing behind Sora. Sora leads with a VideoTI score of 0.90 and Motion Smoothness of 0.99, underscoring its refined alignment of video with the provided text and the fluidity of motion within its generated videos. Mora’s VideoTI score of 0.88 and Motion Smoothness of 0.97, while marginally lower, reflect its robust potential in the accurate interpretation of textual prompts and the generation of smooth motion sequences. Both models exhibit a matched score in Dynamic Degree at 0.75, indicating a comparable proficiency in animating videos with a sense of activity and movement. The Imaging Quality scores show Sora with a slight lead at 0.63 against Mora’s 0.60, suggesting Sora’s subtle superiority in image rendering within video sequences.
Table 3: Text-condtional image-to-video generation. This task evaluates the model’s ability to convert static images and accompanying textual instructions into coherent video sequences. We used 4 pairs of image and text inputs sourced from the official Sora technical report OpenAI (2024b).
Model VideoTI Motion Smoothness Dynamic Dgree Imaging Quality
Sora 0.90 0.99 0.75 0.63
Mora 0.88 0.97 0.75 0.60
Figure 3: Samples for text-conditional image-to-video generation of Mora and Sora. Prompt for the first line image is: Monster Illustration in flat design style of a diverse family of monsters. The group includes a furry brown monster, a sleek black monster with antennas, a spotted green monster, and a tiny polka-dotted monster, all interacting in a playful environment. The second image’s prompt is: An image of a realistic cloud that spells "SORA".
In Figure 3, a qualitative comparison between the video outputs from Sora and Mora reveals that both models adeptly incorporate elements from the input prompt and image. The monster illustration and the cloud spelling "SORA" are well-preserved and dynamically translated into video by both models. Despite quantitative differences, the qualitative results of Mora nearly rival those of Sora, with both models are able to animate the static imagery and narrative elements of the text descriptions into coherent video. This qualitative observation attests to Mora’s capacity to generate videos that closely parallel Sora’s output, achieving a high level of performance in rendering text-conditional imagery into video format while maintaining the thematic and aesthetic essence of the original inputs.
Extend Generated Videos.  The quantitative results from Table 4 reveal that while Sora holds a slight edge over Mora in TCON and Imaging Quality, indicating a higher consistency and fidelity in extending video sequences, Mora’s performance remains close, with a marginal difference of 0.05 in TCON and 0.04 in Imaging Quality. The closeness in Temporal Style scores, 0.24 for Sora and 0.22 for Mora, further signifies that Mora nearly matches Sora in maintaining stylistic continuity over time. Despite Sora’s lead, Mora’s capabilities, particularly in following the temporal style and extending existing videos without significant quality loss, demonstrate its effectiveness in the video extension domain.
Table 4: Extend generated videos. This experiment compares the continuity and quality of the extended video with the same input video given to Sora and Mora respectively. The input video is taken from the full video of the offical Sora technical report OpenAI (2024b), and we take 2 clips as the input video for this task.
Model TCON Imaging Quality Temporal Style
Sora 0.99 0.43 0.24
Mora 0.94 0.39 0.22
Figure 4: Samples for Extend generated video of Mora and Sora.
From a qualitative standpoint, Figure 4 illustrates the competencies of Mora in extending video sequences. Both Sora and Mora adeptly maintain the narrative flow and visual continuity from the original to the extended video. Despite the slight numerical differences highlighted in the quantitative analysis, the qualitative outputs suggest that Mora’s extended videos preserve the essence of the original content with high fidelity. The preservation of dynamic elements such as the rider’s motion and the surrounding environment’s blur effect in the Mora generated sequences showcases its capacity to produce extended videos that are not only coherent but also retain the original’s motion and energy characteristics. This visual assessment underscores Mora’s proficiency in generating extended video content that closely mirrors the original, maintaining the narrative context and visual integrity, thus providing near parity with Sora’s performance.
Table 5: Video-to-video editing. This task examines the model’s ability to edit and transform existing video content in accordance with textual instructions while maintaining visual and stylistic coherence. Data includes 1 input video and 12 instructions from official Sora technical report.
Model Imaging Quality Temporal Style
Sora 0.52 0.24
Mora 0.38 0.23
Video-to-Video Editing.  Table 5 exhibits a comparative analysis of video-to-video editing capabilities between Sora and Mora. Sora secures a higher score in Imaging Quality at 0.52, which suggests a superior capability in preserving the visual details and overall image fidelity during the video editing process. Mora demonstrates capability in video-to-video editing with an Imaging Quality score of 0.38. Although this score reveals a noticeable discrepancy when compare to Sora, it offers valuable insights on pinpointing areas for targeted enhancements in future iterations of the Mora. In the aspect of Temporal Style, both models exhibit proximal performance, with Sora marginally leading at 0.24 compared to Mora’s 0.23. This near-parity underscores the capacity of Mora to closely emulate the stylistic temporal consistency achieved by Sora, ensuring that the edited videos retain a coherent style over time, a crucial aspect of seamless video-to-video editing.
Figure 5: Samples for Video-to-video editing
Table 6: Connect Videos. This task assess the capability of models to seamlessly integrate distinct video clips into a cohesive sequence. $*{}^{*}$ora data we utilize is 5 distinct video-video pairs from Sora technical report website and Mora data we utilize 5 video-video pairs of similar styles.
Model Imaging Quality Tmean
$*{}^{*}$ 0.52 0.64
Mora $*{}^{*}$ 0.42 0.45
Figure 6: Samples for Connect Videos
Upon qualitative evaluation, Figure 5 presents samples from video-to-video editing tasks, wherein both Sora and Mora were instructed to modify the setting to the 1920s style while maintaining the car’s red color. Visually, Sora’s output exhibits a transformation that convincingly alters the modern-day setting into one reminiscent of the 1920s, while carefully preserving the red color of the car. Mora’s transformation, while achieving the task instruction, reveals differences in the execution of the environmental modification, with the sampled frame from generated video suggesting a potential for further enhancement to achieve the visual authenticity displayed by Sora. Nevertheless, Mora ’s adherence to the specified red color of the car underline its ability to follow detailed instructions and enact considerable changes in the video content. This capability, although not as refined as Sora’s, demonstrates Mora’s potential for significant video editing tasks.
Connect Videos.  Quantitative analysis based on the results shown in Table 6 suggests that the model Sora outperforms Mora in terms of Imaging Quality and Tmean. Sora achieves a score of 0.52 in Imaging Quality, indicating its higher fidelity in visual representation compared to Mora’s 0.42. Furthermore, Sora’s superiority is evident in the Temporal coherence aspect with a score of 0.64, which implies that Sora maintains a more consistent visual narrative over time than Mora, which scores 0.45. This quantitative assessment not only solidifies Sora’s position as a superior model in creating high-quality, temporally coherent video sequences but also delineates a trajectory for future enhancements in video connectivity for the Mora framework.
Qualitative analysis based on Figure 6 suggest that, in comparison to Sora’s proficiency in synthesizing intermediate video segments that successfully incorporate background elements from preceding footage and distinct objects from subsequent frames within a single frame, the Mora model demonstrates a blurred background in the intermediate videos, which results in indistinguishable object recognition. Accordingly, this emphasizes the potential for advancing the fidelity of images within the generated intermediate videos as well as enhancing the consistency with the entire video sequence. This would contribute to refining the video connecting process and improving the integration quality of Mora’s model outputs.
Table 7: Simulate digital worlds. This task assesses model’s effectiveness in creating videos that emulate digital or virtual environments, with a focus on preserving the visual details and the distinct appearance style of such worlds. By referring to the video generated by Sora from the official technical report OpenAI (2024b), we used GPT-4 OpenAI (2023b) to build the input prompt for Mora, which contains the "minecraft scene" and the "player character".
Model Imaging Quality Appearance style
Sora 0.62 0.23
Mora 0.52 0.23
Figure 7: Samples for Simulate digital worlds
Simulate Digital Worlds.  In the assessment of digital world simulation capabilities, Table 7 presents a comparative metric analysis between Sora and Mora. Sora exhibits a lead in Imaging Quality with a score of 0.62, indicative of its refined capability to render digital worlds with a higher degree of visual realism and fidelity. Mora, with a score of 0.52, although demonstrating a competent level of performance, falls behind Sora, suggesting areas for improvement in achieving the same level of image clarity and detail. However, both models achieve identical scores in Appearance Style, at 0.23, which reflects a shared ability to adhere to the stylistic parameters of the digital worlds being simulated. This suggests that while there is a difference in the imaging quality, the stylistic translation of textual descriptions into visual aesthetics is accomplished with equivalent proficiency by both models.
Upon qualitative evaluation, Figure 7 presents samples from Simulate digital worlds tasks, wherein both Sora and Mora were instructed to generated video of "Minecraft" scenes. In the top row of frames generated by Sora, we see that the videos maintain high fidelity to the textures and elements typical of digital world aesthetics, characterized by crisp edges, vibrant colors, and clear object definition. The pig and the surrounding environment appear to adhere closely to the style one would expect from a high-resolution game or a digital simulation. These are crucial aspects of performance for Sora, indicating a high-quality synthesis that aligns well with user input while preserving visual consistency and digital authenticity. The bottom row of frames generated by Mora suggests a step towards achieving the digital simulation quality of Sora but with notable differences. Although Mora seems to emulate the digital world’s theme effectively, there is a visible gap in visual fidelity. The images generated by Mora exhibit a slightly muted color palette, less distinct object edges, and a seemingly lower resolution compared to Sora’s output. This suggests that Mora is still in a developmental phase, with its generative capabilities requiring further refinement to reach the performance level of Sora.
5 Discussion
5.1 Strengths of Mora
Innovative Framework and Flexibility.  Mora introduces a groundbreaking multi-agent framework for video generation, significantly advancing the field by enabling a vast array of tasks. This innovative approach not only facilitates text-to-video conversion but also supports the simulation of digital worlds, showcasing unparalleled flexibility and efficiency. Unlike its closed-source counterparts such as Sora, Mora’s open framework design stands out, offering the ability to seamlessly integrate various models. This adaptability allows for the completion of an expanded range of tasks and projects, making Mora an indispensable tool for diverse application scenarios.
Open-Source Contribution.  Mora’s open-source nature is highlighted as a significant contribution to the AI community, encouraging further development and refinement by providing a solid foundation upon which future research can build. This open-source approach not only democratizes access to advanced video generation technologies but also fosters collaboration and innovation in the field. The discussion suggests new ways for future research, including improving the framework’s efficiency, reducing its computational demands during training, and exploring new agent configurations to enhance performance.
5.2 Limitations of Mora
Video Dataset is all you need.  Collecting high-quality video datasets poses significant challenges, primarily due to copyright restrictions on many videos. Unlike images, which can often be easier to gather and utilize for training purposes due to a broader range of available resources and more lenient copyright laws, videos frequently contain copyrighted materials that are not as straightforward to collect or use legally. This issue is compounded by the fact that high-quality videos often come from professional sources, such as movies, television shows, and proprietary game footage, which are rigorously protected by copyright laws. Especially in scenarios involving humans in realistic settings, Mora struggles to generate lifelike movements, such as walking or riding a bicycle. Capturing the subtleties of human motion requires not just any video footage, but high-resolution, smoothly captured sequences that detail every aspect of human kinetics, including the nuances of balance, posture, and interaction with surroundings. Without access to extensive datasets that accurately represent this wide range of human movements, it becomes challenging for video generation models like Mora to replicate these actions convincingly. This limitation underscores the importance of not only the quantity but also the quality and diversity of video datasets in training models to understand and recreate complex human behaviors accurately.
Quality and Length Gaps.  Despite its innovative approach, Mora faces notable challenges, especially when compared to its counterpart, Sora, in terms of video generation quality and capabilities. While Mora is capable of accomplishing tasks similar to those of Sora, the quality of videos generated by Mora falls significantly short, particularly in scenarios involving substantial object movement. Such conditions often introduce a considerable amount of noise, with the quality degradation becoming more pronounced relative to the video length. Moreover, although Mora generally maintains quality for up to 12 seconds of video, any attempt to extend beyond this duration leads to a marked decline in video quality. In contrast, Sora has demonstrated the ability to produce high-quality videos exceeding one minute in length, as showcased in its technical report. This disparity in imaging quality, alongside the limitations in generating videos longer than 12 seconds at a resolution of 1024x576, underscores the urgent need for advancements in Mora’s rendering capabilities and an extension of its video generation parameters.
Instruction Following Capability.  Mora has reached a milestone in its development, now capable of generating videos exceeding 10 seconds from given prompts. However, despite its ability to include all objects specified in the prompts within the generated videos, Mora encounters limitations in executing certain functions. Notably, it struggles with interpreting and rendering the dynamics of motion described in prompts, such as the speed of movement. Additionally, Mora lacks the capability to control the direction of motion—specific directions like left or right for specified objects remain unachievable. This shortfall primarily stems from the system’s foundational approach to video generation, which operates on an Image to Video basis without direct input from textual prompts. Consequently, meeting these specific requirements poses significant challenges, highlighting areas for potential enhancement in Mora’s functionality and interpretative capabilities.
Human Visual Preference Alignment.  The absence of human labeling information within the video domain suggests that experimental results may not always align with human visual preferences, highlighting a significant gap. For example, during connection video task, the generation of transition videos that disintegrate a male figure only to form a female figure represents an exceedingly illogical scenario. This example underscores the necessity for datasets that adhere more closely to physical laws, thereby refining our work and ensuring more realistic and coherent video generation outcomes. Such a framework would enhance the appeal of generated videos to viewers and contribute to establishing comprehensive benchmarks and evaluation criteria for the field.
6 Conclusion
We introduce Mora, a groundbreaking generalist framework for video generation that addresses a range of video-related tasks. Leveraging the collaborative power of multiple agents, Mora marks a considerable advancement in generating videos from textual prompts, establishing new benchmarks for adaptability, efficiency, and output quality in the field of video generation. Our thorough evaluation reveals that Mora not only competes with but also exceeds the capabilities of current leading models in certain areas. However, it has a notable gap with Sora model by OpenAI whose closed-source nature poses considerable challenges for replication and innovation within the academic and professional communities. Our work demonstrates the untapped potential of a meta-programming approach that facilitates intricate collaborations among a variety of agents, each specializing in a segment of the video generation process. In essence, the achievements of Mora not only illustrate the current open source state-of-the-art models in video generation but also illuminate the path forward for the field. As we continue to explore the vast landscape of generative AI, multi-agent collaboration frameworks like Mora will undoubtedly play a pivotal role in unlocking new creative possibilities and applications, from storytelling and content creation to simulation and training. The journey of innovation is far from over, and Mora represents a significant milestone on this ongoing voyage of discovery.
Looking ahead, there are several promising pathways for further research. One such direction involves exploring the integration of more sophisticated natural language understanding capabilities within the agents, potentially allowing for more detailed and context-aware video generations. In addition, the expansion of Mora to incorporate real-time feedback loops could offer interactive video creation experiences, where user inputs could guide the generation process in more dynamic and responsive ways. Furthermore, the challenge of accessibility and computational resource requirements remains a critical barrier to wider adoption and innovation. Future iterations of Mora could benefit from optimizations that reduce these requirements, making advanced video generation technologies more accessible to a broader range of users and developers. In parallel, efforts to create more open and collaborative research environments could accelerate progress in this domain, enabling the community to build upon the foundation laid by Mora framework and other pioneering works.